Book Depot Inc. sells on terms of 3/10, net 60. What is the implicit cost of trade credit under these terms? Use a 365-day year.
Answer #1

IMPLICIT COST OF CREDIT = Discount %/(1-Discount %) x (365/(Full allowed payment days - Discount days))

IMPLICIT COST OF CREDIT = 0.03/(1-0.03) X (365/(60-10))

IMPLICIT COST OF CREDIT = (0.03/0.97) X (365/50) = 22.577%

GO THROUGH IT. ANY DOUBTS, FEEL FREE TO ASK. GIVE POSITIVE FEEDBACK